### v1

@@ Abstract.
We prove that two topologically conjugate bi-critical circle maps whose signatures are the same, and whose renormalizations converge together exponentially fast in the $C^{1}$-topology, are $C^{1}$ conjugate.
@@ 1. Introduction
Rigidity results have also been shown for classes of circle maps with critical points, in particular for multi-critical circle maps. Multi-critical circle maps are $C^{3}$ maps of the circle with a finite number of non-flat critical points. Any non-flat critical point has an associated real number which is called the criticality of the critical point. Yoccoz proved in [Yoc84] that any multi-critical circle map is topologically conjugate to the corresponding rigid rotation, in particular two multi-critical circle maps with the same irrational rotation number are topologically conjugate to each other. We are interested in studying the smoothness of the conjugacy between two critical circle maps. as opposed to the conjugacy between a critical circle map and the corresponding rigid rotation. If each of the maps only have one critical point and the criticalities of the two critical points are the same then the following smoothness results have been shown for the conjugacy $h$ that sends one critical point to the other: $h$ is quasi-symmetric (and therefore Hölder) [dFdM99]; $h$ is a $C^{1}$ diffeomorphism [KT07, GMdM18] and it is a $C^{1+\alpha}$ diffeomorphism for a full Lebesgue measure subset $\mathcal{A}$ of irrational rotation numbers [dFdM99], provided that the critical maps are $C^{4}$. Moreover, it has been shown [Avi13, dFdM99] that there exist real-analytic and $C^{\infty}$ uni-critical circle maps whose rotation numbers are not contained in $\mathcal{A}$ where the conjugacy is only $C^{1}$ and not $C^{1+\alpha}$.
@@
Let $f$ and $g$ be two $C^{3}$ bi-critical circle maps with the same signature. If the renormalizations of $f$ and $g$ around corresponding critical points converge together exponentially fast in the $C^{1}$ topology, then $f$ and $g$ are $C^{1}$-conjugated.
In [EG23], it was shown (under the same condition on the renormalization as in Theorem A) that two $C^{3}$ bi-critical circle maps are $C^{1+\alpha}$-conjugated if their rotation number belongs to a Lebesgue measure total set $\mathcal{A}$. In particular, those bi-critical circle maps are $C^{1}$-conjugate. Therefore, the main contribution in this paper is to dispose of the hypothesis on the rotation number.
@@ 2.1. Multicritical circle maps
A multicritical circle map is an orientation preserving $C^{3}$ circle homeomorphism having $N\geq 1$ non-flat critical points and with irrational rotation number. A critical point $c$ of a $C^{3}$ map $f$ is non-flat of criticality $d>1$ if there exists a neighbourhood $W$ of the critical point such that  $f(x)=f(c)+\phi(x)\,\big{|}\phi(x)\big{|}^{d-1}$ for all $x\in W$, where $\phi:W\rightarrow\phi(W)$ is an orientation preserving $C^{3}$ diffeomorphism satisfying $\phi(c)=0$. As we mentioned in the introduction, any multicritical circle map $f$ has no wandering intervals and in particular it is topologically conjugate with the corresponding rigid rotation [Yoc84].
@@ 2.2. Combinatorics of multicritical circle maps
Truncating the continued fraction expansion, we obtain the so-called convergents of $\rho$, given by $\frac{p_{n}}{q_{n}}=\;[a_{0},a_{1},\cdots,a_{n-1}].$ The sequence of denominators $\{q_{n}\}_{n\in\mathbb{N}}$ satisfies the following recursive formula:
@@
Given $f$ a circle homeomorphism with irrational rotation number given by $\rho$ and given any $x\in S^{1}$, the sequence of iterates $\{f^{q_{n}}(x)\}_{n\in\mathbb{N}}$ converges to $x$ (but never coincides with $x$ since we do not have periodic points) alternating the sides where it appoaches. Moreover, the iterate $f^{q_{n}+q_{n+1}}(x)$ is closer to $x$ than $f^{q_{n}}(x)$ but further to $x$ than $f^{q_{n+1}}(x)$. By orientation, $f^{q_{n}+q_{n+1}}(x)$ must be inside the smallest arc that connects $x$ with $f^{q_{n}}(x)$. We denote by $I_{n}(x)$ the interval with endpoints $x$ and $f^{q_{n}}(x)$, which contains the point $f^{q_{n+2}}(x)$. The collection of intervals
@@
is a partition of the circle (modulo endpoints) called the $n$-th (classical) dynamical partition associated to $x$. The sequence of dynamical partitions $\{\mathcal{P}_{n}(x)\}_{n\in\mathbb{N}}$, for any $x\in S^{1}$, is nested: each element of $\mathcal{P}_{n+1}(x)$ is contained in an element of $\mathcal{P}_{n}(x)$. Those partitions are also refining: the maximal lenght of an element of $\mathcal{P}_{n}(x)$ tends to zero as $n$ tends to infinite. For any $x\in S^{1}$, we denote by $J_{n}(x)$ the union $I_{n}(x)\cup I_{n+1}(x)$ and for any $n\in\mathbb{N}$, we denote by $E_{n}$ the set of endpoints of $\mathcal{P}_{n}(c)$ for $c\in Crit(f)$. With dynamical partitions at hand, we are able to define the renormalization of a critical circle map.
@@ 2.3. Renormalization of bi-critical circle maps
A $C^{3}$ critical commuting pair with two critical points, also called a bi-critical commuting pair, is a pair $\zeta=(\eta,\xi)$ consisting of two $C^{3}$ orientation preserving homeomorphisms $\xi:I_{\xi}\rightarrow\xi(I_{\xi})$ and $\eta:I_{\eta}\rightarrow\eta(I_{\eta})$, satisfying:
@@
Let $\zeta_{1}=(\eta_{1},\xi_{1})$ and $\zeta_{2}=(\eta_{2},\xi_{2})$ be two $C^{r}$ bi-critical commuting pairs, and let $\tau_{1}:[\eta_{1}(0),\xi_{1}(0)]\rightarrow[-1,1]$ and $\tau_{2}:[\eta_{2}(0),\xi_{2}(0)]\rightarrow[-1,1]$ be the two Mœbius transformations given by
@@
Let us recall how we obtain a bi-critical commuting pair from a given bi-critical circle map. Let $f$ be a $C^{r}$ bi-critical circle map with $\rho(f)=\rho\in\mathbb{R}\setminus\mathbb{Q}$ and critical points $c_{0},c_{1}$. For a given $c_{i}$, let $\widehat{f}$ be the lift of $f$ (under the universal covering $t\mapsto c_{i}\cdot\exp(2\pi it)$) such that $10<\widehat{f}(0)<$ 1 (note that $D\widehat{f}(0)=0$). For $n\geq 1$, let $\widehat{I}_{n}(c_{i})$ be the closed interval in $\mathbb{R}$, having the origin as one of its extreme points, which is projected onto $I_{n}(c_{i})$. We define $\xi:\widehat{I}_{n+1}(c_{i})\rightarrow\mathbb{R}$ and $\eta:\widehat{I}_{n}(c_{i})\rightarrow\mathbb{R}$ by $\xi=T^{-p_{n}}\circ\widehat{f}^{q_{n}}$ and $\eta=T^{-p_{n+1}}\circ\widehat{f}^{q_{n+1}}$, where $T(x)=x+1$. Then the pair $(\eta|_{\widehat{I}_{n}(c_{i})},\xi|_{\widehat{I}_{n+1}(c_{i})})$ is a renormalizable bi-critical commuting pair, and its rescaling is denoted by $\mathcal{R}_{i}^{n}f=\left(f^{q_{n+1}}|_{I_{n}(c_{i})},f^{q_{n}}|_{I_{n+1}(c_{i})}\right).$ In other words, let us consider the $n-$th scaling ratio of $f$:
@@
As a straightforward consequence of the combinatorics, we know that $f$ has irrational rotation number if and only if $f$ is infinitely renormalizable. In fact, once we have the $n-$th renormalization of a pair we can obtain the $(n+1)-$th renormalization by iterating $\mathcal{R}_{i}^{n}f(B_{n,f}(1))$ as many times until we observe a change in the signal of the images. By combinatorics that number coincides with $\ a_{n+1}+1$, since
@@ 2.4. Real bounds
In other words, real bounds say that the lenghts of adjacent intervals in the dynamical partition are comparable. Real bounds are also true even if the circle map has not only two but a finite number of critical points (see [EdF18, EdFG18]).
The following result is a consequence of universal real bounds.
@@ 2.5. Two tools: Schwarzian derivative and Koebe Distortion Principle
Next result give us a decomposition of the iterates of intervals in the dyamical partition.
@@
    For the reminer values of $i$, $\widehat{\phi}_{i}$ is either the identity or a diffeomorphism with negative Schwarzian derivative.
@@ 2.6. Controlled bi-critical commuting pairs
Real bounds also give us some control on the lenght of certain intervals, dynamically defined, and some control on the first and second derivatives of the renormalizations.
@@
($M$-controlled bi-critical communting pair) Let $M>1$ and let $\zeta=(\eta,\xi)$ be a normalized $C^{3}$ bi-critical commuting pair, with period $a=\chi(\zeta)\in\mathbb{N}$ and with critical points given by 00 and $\beta$. We say that $\zeta$ is $M$-controlled if the following conditions are satisfied:
@@
Next result guarantes that every $C^{3}$ bi-critical circle map with irrational rotation number gives rise to a sequence of bi-critical controlled commuting pairs, after a finite number of renormalizations.
@@
By universal real bounds (Theorem 2.1), there exists a constant $C>1$ and $n_{0}\in\mathbb{N}$ such that for any $n\geq n_{0}$ properties $(1)$ to $(5)$ and property $(8)$ hold. Moreover, by [EdF18, Lemma 4.2, Lemma 4.1] and real bounds we know that $\left|[\eta(\beta),\beta]\right|$ is comparable with $\xi(0)$, then by Property $(1)$ $|\eta(\beta)-\beta|\geq 1/M$. By [EdF18, Lemma 3.3], $\left|\beta-\eta^{-1}(\beta)\right|\geq 1/M$, and hence property $(9)$ is also true. Finally, from [EdFG18, Lemma 5.1] and [dFdM99, Lemma A.4], there exists $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ properties $(6)$ and $(7)$ hold. ∎
@@
By [dFdM99, Lemma A4], the renormalizations of a $C^{3}$ uni-critical circle map are controlled as in our definition. That result and its proof can be generalized to the bi-critical case.
@@ 3. The two-bridges partition $\widehat{\mathcal{P}}_{n}$
In this section we recall the construct of the sequence of two-bridges partitions $\{\widehat{\mathcal{P}}_{n}\}_{n\in\mathbb{N}}$, defined in [EG23, Section 4.1.1]. Let $c_{0}$ and $c_{1}$ be the two critical points of $f$. For $n\in\mathbb{N}$, the first return map to $J_{n}(c_{0})$, which is given by $(f^{q_{n+1}}|_{I_{n}(c_{0})},f^{q_{n}}|_{I_{n+1}(c_{0})})$, has two critical points: one being $c_{0}$ and the other one being the unique preimage of $c_{1}$ for the return. Such preimage in $J_{n}(c_{0})$ is called the free critical point at level $n$, and is denoted by $\mathfrak{c}_{n,0}$. Without lost of generality, let us assume that $\mathfrak{c}_{n,0}$ belongs to $I_{n}(c_{0})$. This means that there exists $j<q_{n+1}$ such that $c_{1}\in f^{j}(I_{n}(c_{0}))$ and $f^{j}(\mathfrak{c}_{n,0})=c_{1}$. Note that if $c_{1}$ belongs to the positive orbit of $c_{0}$ then there exists $m\in\mathbb{N}$ such that for all $n\geq m$ the return maps $(f^{q_{n+1}}|_{I_{n}(c_{0})},f^{q_{n}}|_{I_{n+1}(c_{0})})$ are unicritical and we will be in the situation studied in [KT07]. So, from now on, let us assume that $c_{1}$ does not belong to the positive orbit of $c_{0}$. Moreover we will assume that $n\geq n_{0}$ where $n_{0}$ is given by Theorem 2.6.
@@
  • If $v$ is the middle point of $I^{\prime}$, then we add $v$ to the set of endpoints of $\widehat{\mathcal{P}}_{n+1}$.

### v2

@@ Abstract.
We prove that two topologically conjugate bi-critical circle maps whose signatures are the same, and whose renormalizations converge together exponentially fast in the $C^{2}$-topology, are $C^{1}$ conjugate.
@@ 1. Introduction
Rigidity results have also been shown for classes of circle maps with critical points, in particular for multi-critical circle maps. Multi-critical circle maps are $C^{3}$ maps of the circle with a finite number of non-flat critical points. Any non-flat critical point has an associated real number which is called the criticality of the critical point. Yoccoz proved in [Yoc84] that any multi-critical circle map is topologically conjugate to the corresponding rigid rotation, in particular two multi-critical circle maps with the same irrational rotation number are topologically conjugate to each other. We are interested in studying the smoothness of the conjugacy between two critical circle maps as opposed to the conjugacy between a critical circle map and the corresponding rigid rotation. If each of the maps only have one critical point and the criticalities of the two critical points are the same then the following smoothness results have been shown for the conjugacy $h$ that sends one critical point to the other: $h$ is quasi-symmetric (and therefore Hölder) [dFdM99]; $h$ is a $C^{1}$ diffeomorphism [KT07, GMdM18] and it is a $C^{1+\alpha}$ diffeomorphism for a full Lebesgue measure subset $\mathcal{A}$ of irrational rotation numbers [dFdM99], provided that the critical maps are $C^{4}$. Moreover, it has been shown [Avi13, dFdM99] that there exist real-analytic and $C^{\infty}$ uni-critical circle maps whose rotation numbers are not contained in $\mathcal{A}$ where the conjugacy is only $C^{1}$ and not $C^{1+\alpha}$.
@@
Let $f$ and $g$ be two $C^{3}$ bi-critical circle maps with the same signature. If the renormalizations of $f$ and $g$ around corresponding critical points converge together exponentially fast in the $C^{2}$ topology, then $f$ and $g$ are $C^{1}$-conjugated.
In [EG23], it was shown that two $C^{3}$ bi-critical circle maps $f,g$ are $C^{1+\alpha}$-conjugate if their rotation number belongs to a total Lebesgue measure set $\mathcal{A}$, provided that their renormalizations converge together in the $C^{1}$ topology (instead of the $C^{2}$ topology required in Theorem A). In particular those bi-critical circle maps are $C^{1}$-conjugate. The condition on the rotation number belonging to $\mathcal{A}$ let the authors avoid the use of the $C^{2}$ topology. In this paper we are not imposing any condition on the irrational rotation number, due to this we have to deal with the existence of almost tangency points as well as their first two derivatives under renormalization (see Section 5). Assuming that the renormalizations of $f$ and $g$ converge together exponentially in the $C^{2}$ topology allows us to obtain a suitable control on the first two derivatives of tangency points of $f$ and $g$.
@@ 2.1. Multicritical circle maps
A multicritical circle map is an orientation preserving $C^{3}$ circle homeomorphism having $N\geq 1$ non-flat critical points and with irrational rotation number. A critical point $c$ of a $C^{3}$ map $f$ is non-flat of criticality $d>1$ if there exists a neighborhood $W$ of the critical point such that $f(x)=f(c)+\phi(x)\,\big{|}\phi(x)\big{|}^{d-1}$ for all $x\in W$, where $\phi:W\rightarrow\phi(W)$ is an orientation preserving $C^{3}$ diffeomorphism satisfying $\phi(c)=0$. As we mentioned in the introduction, any multicritical circle map $f$ has no wandering intervals and in particular it is topologically conjugate with the corresponding rigid rotation [Yoc84].
@@ 2.2. Combinatorics of multicritical circle maps
Truncating the continued fraction expansion, we obtain the convergents of $\rho$, given by $\frac{p_{n}}{q_{n}}=\;[a_{0},a_{1},\cdots,a_{n-1}].$ The sequence of denominators $\{q_{n}\}_{n\in\mathbb{N}}$ satisfies the following recursive formula:
@@
Given a preserving orientation circle homeomorphism $f$ with irrational rotation number $\rho$ and given any $x\in S^{1}$, the sequence of iterates $\{f^{q_{n}}(x)\}_{n\in\mathbb{N}}$ converges to $x$ (but never coincides with $x$ since $f$ has no periodic points) alternating the sides where it approaches. Moreover, the iterate $f^{q_{n}+q_{n+1}}(x)$ is closer to $x$ than $f^{q_{n}}(x)$ but further to $x$ than $f^{q_{n+1}}(x)$. By orientation, $f^{q_{n}+q_{n+1}}(x)$ must be inside the smallest arc that connects $x$ with $f^{q_{n}}(x)$. We denote by $I_{n}(x)$ the interval with endpoints $x$ and $f^{q_{n}}(x)$, which contains the point $f^{q_{n+2}}(x)$. The collection of intervals
@@
is a partition of the circle (modulo endpoints) called the $n$-th (classical) dynamical partition associated to $x$. The sequence of dynamical partitions $\{\mathcal{P}_{n}(x)\}_{n\in\mathbb{N}}$, for any $x\in S^{1}$, is nested: each element of $\mathcal{P}_{n+1}(x)$ is contained in an element of $\mathcal{P}_{n}(x)$. Those partitions are also refining: the maximal length of an element of $\mathcal{P}_{n}(x)$ converges to zero. For any $x\in S^{1}$, we denote by $J_{n}(x)$ the union $I_{n}(x)\cup I_{n+1}(x)$ and for any $n\in\mathbb{N}$, we denote by $E_{n}$ the set of endpoints of $\mathcal{P}_{n}(c)$ for $c\in Crit(f)$. With dynamical partitions at hand, we are able to define the renormalization of a critical circle map.
@@ 2.3. Renormalization of bi-critical circle maps
A $C^{3}$ critical commuting pair with two critical points, or a bi-critical commuting pair, is a pair $\zeta=(\eta,\xi)$ consisting of two $C^{3}$ orientation preserving homeomorphisms $\xi:I_{\xi}\rightarrow\xi(I_{\xi})$ and $\eta:I_{\eta}\rightarrow\eta(I_{\eta})$, satisfying:
@@
Let $\zeta_{1}=(\eta_{1},\xi_{1})$ and $\zeta_{2}=(\eta_{2},\xi_{2})$ be two $C^{r}$ bi-critical commuting pairs, and let $\tau_{1}:[\eta_{1}(0),\xi_{1}(0)]\rightarrow[-1,1]$ and $\tau_{2}:[\eta_{2}(0),\xi_{2}(0)]\rightarrow[-1,1]$ be the two Mœbius transformations with
@@
Let us recall how we obtain a bi-critical commuting pair from a given bi-critical circle map. Let $f$ be a $C^{r}$ bi-critical circle map with $\rho(f)=\rho\in\mathbb{R}\setminus\mathbb{Q}$ and critical points $c_{0},c_{1}$. For a given $c_{i}$, let $\widehat{f}$ be the lift of $f$ (under the universal covering $t\mapsto c_{i}\cdot\exp(2\pi it)$) such that $10<\widehat{f}(0)<$ 1 (note that $D\widehat{f}(0)=0$). For $n\geq 1$, let $\widehat{I}_{n}(c_{i})$ be the closed interval in $\mathbb{R}$, that has the origin as one of its endpoints and which projects onto $I_{n}(c_{i})$. We define $\xi:\widehat{I}_{n+1}(c_{i})\rightarrow\mathbb{R}$ and $\eta:\widehat{I}_{n}(c_{i})\rightarrow\mathbb{R}$ by $\xi=T^{-p_{n}}\circ\widehat{f}^{q_{n}}$ and $\eta=T^{-p_{n+1}}\circ\widehat{f}^{q_{n+1}}$, where $T(x)=x+1$. Then the pair $(\eta|_{\widehat{I}_{n}(c_{i})},\xi|_{\widehat{I}_{n+1}(c_{i})})$ is a renormalizable bi-critical commuting pair, and its rescaling is denoted by $\mathcal{R}_{i}^{n}f=\left(f^{q_{n+1}}|_{I_{n}(c_{i})},f^{q_{n}}|_{I_{n+1}(c_{i})}\right).$ In other words, let us consider the $n-$th scaling ratio of $f$:
@@
As a straightforward consequence of the combinatorics, we know that $f$ has irrational rotation number if and only if $f$ is infinitely renormalizable. In fact, once we have the $n-$th renormalization of a pair we can obtain the $(n+1)-$th renormalization by iterating $\mathcal{R}_{i}^{n}f(B_{n,f}(1))$ as many times until we observe a change in the sign of the images. By combinatorics that number coincides with $\ a_{n+1}+1$, since
@@ 2.4. Real bounds
In other words, real bounds say that the lengths of adjacent intervals in the dynamical partition are comparable. Real bounds are also true even if the circle map has not only two but a finite number of critical points (see [EdF18, EdFG18]).
The following result is a consequence of having universal real bounds.
@@ 2.5. Two tools: Schwarzian derivative and Koebe Distortion Principle
Next result give us a decomposition of the iterates of intervals in the dynamical partition.
@@
    For the reminder values of $i$, $\widehat{\phi}_{i}$ is either the identity or a diffeomorphism with negative Schwarzian derivative.
@@ 2.6. Controlled bi-critical commuting pairs
Real bounds also give us some control on the length of certain intervals, dynamically defined, and some control on the first and second derivatives of the renormalizations.
@@
($M$-controlled bi-critical commuting pair) Let $M>1$ and let $\zeta=(\eta,\xi)$ be a normalized $C^{3}$ bi-critical commuting pair, with period $a=\chi(\zeta)\in\mathbb{N}$ and with critical points given by 00 and $\beta\in(0,\xi(0))$. We say that $\zeta$ is $M$-controlled if the following conditions are satisfied:
@@
Analogously, in the case that $\beta\in(\eta(0),0)$, we can extend the definition of $M$-controlled bi-critical commuting pair.
The next result guarantees that every $C^{3}$ bi-critical circle map with irrational rotation number gives rise to a sequence of bi-critical controlled commuting pairs, after a finite number of renormalizations.
@@
By universal real bounds (Theorem 2.1), there exists a constant $C>1$ and $n_{0}\in\mathbb{N}$ such that for any $n\geq n_{0}$ properties $(1)$ to $(5)$ and property $(8)$ hold. Moreover, by [EdF18, Lemma 4.2, Lemma 4.1] and real bounds we know that $\left|[\eta(\beta),\beta]\right|$ is comparable with $\xi(0)$, then by Property $(1)$ $|\eta(\beta)-\beta|\geq 1/M$. By [EdF18, Lemma 3.3], $\left|\beta-\eta^{-1}(\beta)\right|\geq 1/M$, and hence property $(9)$ is also true. Finally, from [EdFG18, Lemma 5.1] and [dFdM99, Theorem A.4 (pages 389-381) and remark (page 381)], there exists $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ properties $(6)$ and $(7)$ hold. ∎
@@
By [dFdM99, Theorem A.4 (pages 389-381) and remark (page 381)], the renormalizations of a $C^{3}$ uni-critical circle map are controlled as in our definition. That result and its proof can be generalized to the bi-critical case.
As $\mathcal{R}^{n}f$ and $\mathcal{R}^{n}g$ satisfy Theorem 2.6 and Inequality (1) for $d=2$, we know that $\lim_{n\to\infty}\log\frac{|h(I_{n}(c_{i}))|}{|I_{n}(c_{i})|}$ exists and converge exponentially. Conjugating the map $f$ (or $g$) by a $C^{\infty}$ diffeomorphism we can achieve $\lim_{n\to\infty}\log\frac{|h(I_{n}(c_{i}))|}{|I_{n}(c_{i})|}=0$ at an exponential rate. Therefore, we can always assume that there exist $C>0$ and $\lambda\in\{0,1\}$ such that for any $n\geq n_{0}$ and $i\in(0,1)$ we have
$\left|\log\frac{|h(I_{n}(c_{i}))|}{|I_{n}(c_{i})|}\right|\leq C\,\lambda^{n}.$ (2)
@@ 3. The two-bridges partition $\widehat{\mathcal{P}}_{n}$
In this section we recall the construct of the sequence of the two-bridges partitions $\{\widehat{\mathcal{P}}_{n}\}_{n\in\mathbb{N}}$, defined in [EG23, Section 4.1.1]. Let $c_{0}$ and $c_{1}$ be the two critical points of $f$. For $n\in\mathbb{N}$, the first return map to $J_{n}(c_{0})$, which is given by $(f^{q_{n+1}}|_{I_{n}(c_{0})},f^{q_{n}}|_{I_{n+1}(c_{0})})$, has two critical points: one being $c_{0}$ and the other one being the unique pre-image of $c_{1}$ in $J_{n}(c_{0})$. Such a preimage in $J_{n}(c_{0})$ is called the free critical point at level $n$, and is denoted by $\mathfrak{c}_{n,0}$. Without lost of generality, let us assume that $\mathfrak{c}_{n,0}$ belongs to $I_{n}(c_{0})$. This means that there exists $j<q_{n+1}$ such that $c_{1}\in f^{j}(I_{n}(c_{0}))$ and $f^{j}(\mathfrak{c}_{n,0})=c_{1}$. Note that if $c_{1}$ belongs to the positive orbit of $c_{0}$ then there exists $m\in\mathbb{N}$ such that for all $n\geq m$ the return maps $(f^{q_{n+1}}|_{I_{n}(c_{0})},f^{q_{n}}|_{I_{n+1}(c_{0})})$ are unicritical and we will be in the situation studied in [KT07]. So, from now on, let us assume that $c_{1}$ does not belong to the positive orbit of $c_{0}$. Moreover we will assume that $n\geq n_{0}$ where $n_{0}$ is given by Theorem 2.6.
@@
  • If $v$ is the mid point of $I^{\prime}$, then we add $v$ to the set of endpoints of $\widehat{\mathcal{P}}_{n+1}$.